# A Plug-and-Play Learning-based IMU Bias Factor for Robust Visual-Inertial Odometry

### v1

@@ Abstract
The bias of low-cost Inertial Measurement Units (IMU) is a critical factor affecting the performance of Visual-Inertial Odometry (VIO). In particular, when visual tracking encounters errors, the optimized bias results may deviate significantly from the true values, adversely impacting the system’s stability and localization precision. In this paper, we propose a novel plug-and-play framework featuring the Inertial Prior Network (IPNet), which is designed to accurately estimate IMU bias. Recognizing the substantial impact of initial bias errors in low-cost inertial devices on system performance, our network directly leverages raw IMU data to estimate the mean bias, eliminating the dependency on historical estimates in traditional recursive predictions and effectively preventing error propagation. Furthermore, we introduce an iterative approach to calculate the mean value of the bias for network training, addressing the lack of bias labels in many visual-inertial datasets. The framework is evaluated on two public datasets and one self-collected dataset. Extensive experiments demonstrate that our method significantly enhances both localization precision and robustness, with the ATE-RMSE metric improving on average by 46%. The source code and video will be available at https://github.com/yiyscut/VIO-IPNet.git.
@@ I INTRODUCTION
In recent years, high-precision and robust localization technologies have become critical in industrial applications, particularly in robotics [1, 2] and augmented reality [3]. As the cost and size of vision sensors and Inertial Measurement Units (IMU) continue to decline, their adoption has become increasingly widespread. The complementary nature of vision sensors and IMU in data measurement has further propelled their integration, establishing Visual-Inertial Odometry (VIO) as a key method for achieving advanced localization capabilities.
The IMU provides high-frequency angular velocity and acceleration data that effectively compensate for tracking errors in vision-only systems. However, the measurements are prone to drift due to factors such as temperature, a challenge particularly pronounced in low-cost devices. This drift is typically modeled as the combination of bias error and random error. When visual feature quality is high, visual data predominates, and the estimator leverages the constraints between visual and inertial data to estimate the bias [4]. Conversely, when visual features are compromised, optimization-based VIO frameworks often lack robust constraints on IMU bias estimation, leading to significant deviations between the estimated and actual bias values over certain time intervals. These issues undermine the precision and robustness of localization systems, limiting their effectiveness in demanding scenarios.
Figure 1: Comparison of IPNet with recursive prediction methods for IMU bias estimation. IPNet directly estimates bias from raw IMU data, preventing error propagation, while recursive methods rely on historical estimates. The prior factors constructed by it significantly enhance both localization precision and robustness.
@@
Figure 2: The overview of the entire framework. Given image and IMU data, the framework outputs real-time pose information. At first, the IMU data is passed through the IPNet to estimate the current bias prior. Subsequently, the data stream is divided into two parts: one part inputs the bias prior along with the raw measurements into the pre-integration module, and performs optimization in the backend using IMU factors; the other part incorporates it as a prior factor into the factor graph, imposing reasonable constraints on the bias optimization process to ensure the precision of the bias estimation.
Existing methods for IMU bias estimation in optimization-based VIO frameworks face significant limitations. For instance, the baseline approach using monocular and inertial data in VINS-Fusion [5] struggles with unfounded adjustments of the bias during optimization, often forcing the bias to fit visual constraints inaccurately. In recent years, researchers have attempted to address this issue using deep learning methods, as seen in [6], by modeling the dynamic evolution of IMU bias. However, the proposed method heavily rely on the estimation results from the previous time step, leading to continuous error propagation once a mistake occurs. This problem is particularly pronounced when the initial values are inaccurate, and it can even lead to trajectory divergence. To overcome these limitations, we propose the Inertial Prior Network (IPNet), which directly estimates IMU bias in real time from raw IMU data, eliminating the dependency on historical estimates in traditional recursive predictions. Furthermore, considering the difficulty of obtaining real-time, accurate IMU bias labels during motion, we use the average bias of each sequence as the label. This approach simplifies the training process while maintaining accuracy. Finally, recognizing that many visual-inertial datasets lack the ground truth bias, we introduce an iterative method to compute the average bias, which is then used for network training and experimentation. Overall, our main contributions are as follows:
@@
(i) We propose IPNet, a novel network for real-time IMU bias estimation from raw data, eliminating recursive prediction dependency and preventing error propagation. By constructing bias estimates as a plug-and-play prior factor integrated into the VIO framework, the approach significantly enhances localization precision and robustness.
(ii) We introduce an iterative method to compute mean IMU bias, addressing the lack of ground truth in public datasets and providing reliable training labels.
@@
(iii) Extensive experiments on public and self-collected datasets demonstrate significant improvements in localization precision and robustness.
@@ II-A Visual Inertial Odometry
The VIO algorithm was initially developed using the Extended Kalman Filter (EKF), primarily due to the computational limitations of early systems. This line of research has continued to evolve, with notable contributions including MSCKF [7], ROVIO [8], MSCKF 2.0 [9], OpenVINS [10], and SchurVINS [11]. With advancements in nonlinear optimization theory and improvements in computational power, optimization-based VIO methods have become the dominant approach. These methods integrate observational data and system constraints by formulating optimization problems, effectively overcoming the limitations of traditional filtering methods in handling complex nonlinear systems. Notable examples include OKVIS [12], VINS-Mono [4], ORB-SLAM3 [13], BASALT [14], Kimera [15], and DM-VIO [16].
However, existing methods often struggle to perform reliably under extreme conditions, such as low-texture environments or fast motion. These challenging scenarios can significantly degrade the accuracy of state estimation or even lead to system failures. As a result, improving the robustness of VIO algorithms remains a critical challenge in the field.
Figure 3: The architecture of IPNet consists of encoder block, sequence modeling block and decoder block. The encoder block is mainly composed of four residual blocks and max-pooling layers, used to extract features and reduce the data dimension. The sequence modeling block is designed to further capture temporal dependencies and the decoder block is responsible for performing regression to predict the bias of both acceleration and angular velocity.
@@ II-B Learning Based Inertial Odometry
In addition to trajectory estimation, many learning-based methods focus on IMU noise reduction and uncertainty estimation. For example, Brossard et al. [22] employed Convolutional Neural Networks (CNN) for gyro denoising. Steinbrener et al. [23] compared different architectures for denoising IMU measurements and found that Long Short-Term Memory Networks (LSTM) outperform Transformer architectures, especially under varying data rates. AirIMU [24] introduced an encoder-decoder network structure to process raw IMU data, estimating IMU states and uncertainties, and outputting corrected acceleration and angular velocity values along with their associated uncertainties. Zeinali et al. [25] proposed MobileResNet, a CNN-based lightweight block that utilizes deep and pointwise convolutions to reduce computational costs and noise interference, thereby improving navigation and localization. Moreover, Guo et al. [26] developed a dead reckoning algorithm for wheeled mobile robots by combining an Invariant Extended Kalman Filter with an IMU error model, enhanced by deep learning to estimate noise parameters. Yang et al. [27] integrated the Left-Invariant Extended Kalman Filter with a statistical neural network, proposing a learning-based dead reckoning algorithm that relies solely on inertial measurements and exhibits robust performance under sudden changes in lighting conditions.
However, existing methods still face challenges due to integration drift caused by accumulated noise and bias, which reduces the long-term accuracy of trajectory estimation. To address this issue, researchers combine IMU data with visual sensors and leverage deep learning to further improve both accuracy and robustness.
@@ II-C Learning Based Visual Inertial Odometry
Unlike existing approaches, we revisit the network model for IMU bias prediction and propose a deep network that directly regresses the mean bias from raw IMU measurements, as detailed in Section III-B. Furthermore, we further developed an IMU prior factor based on this network, which can be easily integrated into any VIO method.
@@ III METHODOLOGY
The overview of our proposed framework is shown in Fig. 2. We use VINS-Fusion [5] as the baseline and integrate the plug-and-play learning-based IMU bias factor into it. This integration addresses a key challenge in VIO frameworks: the degradation of localization precision and robustness caused by inaccurate bias estimation due to visual tracking errors. Furthermore, to facilitate the acquisition of IMU bias labels for network training, we introduce a method for iteratively solving the bias using the ground truth pose and the ground truth velocity.
@@ III-A2 Preintegration
We follow the discrete expressions in [30] and express the displacement, velocity, and rotational preintegration terms as $\boldsymbol{\alpha}_{b_{k+1}}^{b_{k}}$,$\boldsymbol{\beta}_{b_{k+1}}^{b_{k}}$,$\gamma_{b_{k+1}}^{b_{k}}$.
[…]
@@ III-B IPNet Model
We designed IPNet to directly predict the bias from raw IMU data. The network consists of three blocks: encoder block, sequence modeling block and decoder block, as shown in Fig.3. The raw IMU data represent the acceleration and angular velocity measurements in three dimensions, and we take a sample window of length $s$ as the input to the model, so that the measurements can be denoted as $\hat{\mathbf{a}}_{t}\in\mathbb{R}^{s\times 3}$ and $\hat{\boldsymbol{\omega}}_{t}\in\mathbb{R}^{s\times 3}$. We predict $n$ values for each window to avoid prediction fluctuations. Therefore, the model outputs the predicted bias $\hat{\mathbf{b}}_{a}\in\mathbb{R}^{n\times 3}$ and $\hat{\mathbf{b}}_{w}\in\mathbb{R}^{n\times 3}$. In this paper, we set $s$ to 1000 and $n$ to 50.
@@ III-B1 Encoder Block
The encoder block mainly consists of several transformation layers to reshape the tensor and an encoder to capture local features and reduce data dimension. Firstly, concatenate $\hat{\mathbf{a}}_{t}$ and $\hat{\boldsymbol{\omega}}_{t}$ along the last dimension, and then use the transpose operation to swap the positions of the last two dimensions, so as to facilitate the subsequent convolution operations. The encoder contains four sets of residual blocks and max-pooling layers. The main path of residual block primarily utilizes 1D convolution, batch normalization, and PReLU[31] activation, and the residual connection employs a pointwise convolution layer to combine the features of different channels and make the input and output have the same dimension, which are then added to the main path. Finally, a transpose operation is applied to restore the original dimensional order.
@@ III-B3 Decoder Block
To accommodate the characteristics of acceleration bias and angular velocity bias, the decoder block constructs separate accelerometer bias decoder and gyroscope bias decoder using linear layers and transpose operations to predict the bias for the three dimensions of acceleration and angular velocity, respectively. Additionally, to avoid instability caused by fluctuations in model predictions, each decoder outputs $n$ predicted values, and their mean is used as the final prediction for each data window.
@@ III-B4 Loss Function
Figure 4: The comparison of bias values curves corresponding to each pose in the V1_02 sequence. In the baseline, the bias values corresponding to the first 200 poses deviate significantly from the ground truth, whereas with our method, the bias values are quickly adjusted to fluctuate close to the ground truth. Blue: the bias predicted by baseline, Green: the bias predicted by our framework, Red: the ground truth bias.
III-C IMU Bias Prior Factor
The main purpose of this factor is to provide a relatively accurate initial value for the bias optimization. It evaluates the deviation between the actual bias value and the known initial value in the form of residual, and incorporates this constraint into the optimization problem, helping to reduce state drift. The definition of its residual is:
@@ III-C IMU Bias Prior Factor
III-D IMU-Bias Label Solving Method
First, the raw IMU data is pre-integrated to obtain the measured values of the pre-integrated terms. These values are then subtracted from the ground truth of the pre-integrated terms to calculate the error. Finally, $\mathbf{b}_{a_{t}}$ and $\mathbf{b}_{w_{t}}$ are inversely solved through iterative calculations based on the first-order term of the pre-integrated with respect to bias changes. The formula is as follows:

### v2

@@ Abstract
Accurate and reliable estimation of biases of low-cost Inertial Measurement Units (IMU) is a key factor to maintain the resilience of Visual-Inertial Odometry (VIO), particularly when visual tracking fails in challenging areas. In such cases, bias estimates from the VIO can deviate significantly from the real values because of the insufficient or erroneous vision features, compromising both localization accuracy and system stability. To address this challenge, we propose a novel plug-and-play module featuring the Inertial Prior Network (IPNet), which infers an IMU bias prior by implicitly capturing the motion characteristics of specific platforms. The core idea is inspired intuitively by the observation that different platforms exhibit distinctive motion patterns, while the integration of low-cost IMU measurements suffers from unbounded error that quickly accumulates over time. Therefore, these specific motion patterns can be exploited to infer the underlying IMU bias. In this work, we first directly infer the biases prior only using the raw IMU data using a sliding window approach, eliminating the dependency on recursive bias estimation combining visual features, thus effectively preventing error propagation in challenging areas. Moreover, to compensate for the lack of ground-truth bias in most visual-inertial datasets, we further introduce an iterative method to compute the mean per-sequence IMU bias for network training and release it to benefit society. The framework is trained and evaluated separately on two public datasets and a self-collected dataset. Extensive experiments show that our method significantly improves localization precision and robustness. On the EuRoC and TumVi datasets, ATE-RMSE improves by 46% on average, and RPE-RMSE improves by 48% on average. On the In-house dataset(indoor), ATE-RMSE also improves by 46% on average. Furthermore, we apply the model trained in the indoor environment to outdoor scenes and perform testing. The experimental results demonstrate that the model can be successfully applied to outdoor autonomous vehicle. To benefit society, we will open-source our code and release the IMU dataset upon acceptance.
@@ I Introduction
High-precision and robust localization technologies are essential for robotics [1, 2] and augmented reality applications [3]. With the miniaturization and cost reduction of visual sensors and Inertial Measurement Units (IMU), Visual-Inertial Odometry (VIO) has emerged as a leading solution for localization. However, in challenging scenarios where visual data is compromised (such as motion blur, illumination changing, or low-texture environments), the reliability of IMU bias estimation for the low-cost devices becomes critical to VIO performance [4]. More specifically, in the absence of sufficient visual features, IMU bias estimates from the existing VIO systems can deviate significantly from the true values due to the backend optimization of VIO, leading to a degradation in both localization accuracy and system resilience.
Figure 1: Our method ensures that the estimated IMU bias is more physically consistent, particularly when visual errors are significant, thereby mitigating the degradation of localization precision and system robustness.
Existing methods for IMU bias estimation in optimization-based VIO in challenging environments face significant limitations. For instance, the baseline approach using monocular and inertial data in VINS-Fusion [5] struggles with unbounded adjustments of the bias during optimization, often forcing the bias to fit inaccurate visual constraints caused by motion blur or unbalanced features. HCTO [2] attempts to improve the accuracy of low-cost IMU considering human motion patterns, but this method can not be directly applied to other platforms. To address these issues, researchers have turned to deep learning methods to model the dynamic evolution of IMU biases. Similar to our method, HelmetPoser [6] uses a data-driven approach to estimate head motion using IMU data only, but the pure IMU odometry still faces large drift in diverse scenes. Unlike the recursive framework in [7], which estimates time-varying bias sequentially and is therefore sensitive to initialization errors and prone to long-term drift, our non-recursive design infers a stable bias prior directly from raw data. This fundamentally prevents error propagation and improves robustness against label noise by relying on per-sequence mean bias supervision instead of time-varying bias estimation.
@@
Figure 2: The overview of the proposed system. Given image and IMU data, the framework outputs real-time pose information. At first, the IMU data is passed through the IPNet to estimate the current bias prior. Subsequently, the data stream is divided into two parts: one part inputs the bias prior along with the raw measurements into the pre-integration module, and performs optimization in the backend using IMU factors; the other part incorporates it as a prior factor into the factor graph, imposing reasonable prior constraints on the bias optimization process to ensure the robustness of the bias estimation.
From the perspective of traditional inertial navigation theory, the observability of IMU bias strictly relies on external observation information (such as vision, GPS, etc.), and pure IMU data alone cannot provide sufficient observation conditions[8, 9, 10]. However, it is worth noting that certain platform follows specific motion patterns[7], and data-driven neural networks can capture the implicit relationship between IMU bias and motion characteristics, thereby overcoming the limitations of traditional observability frameworks. Therefore, we propose the Inertial Prior Network (IPNet), which learns a stable bias prior—the most probable bias configuration consistent with the platform’s motion dynamics. IPNet infers this prior directly from raw IMU data using a non-recursive architecture, thereby eliminating the accumulation and propagation of historical errors in visually challenging environments. The output serves as a prior factor injected into the VIO backend, where it collaborates with visual observations through a dynamic constraint mechanism for joint optimization, ensuring the physical plausibility of the bias estimates and significantly improving robustness and positioning accuracy. Furthermore, since the backend optimization can compensate for bias estimate residuals, we simplify the training process by using the average bias of each sequence (in visual feature-rich environments) as the supervision signal, and independently train and evaluate on different datasets. Finally, addressing the common issue in visual-inertial datasets of lacking true bias values, we propose an iterative optimization method based on pose ground truth to compute the sequence’s average IMU biases, which is then directly used as the training label for the network. Overall, our main contributions are as follows:
@@
(i) We propose IPNet, a novel network for real-time IMU bias prior inference from raw data, eliminating recursive prediction dependency and preventing error propagation in the visual challenging environment. By constructing bias prior as a plug-and-play prior factor integrated into the VIO framework, the approach significantly enhances localization precision and resilience.
(ii) We introduce an iterative method to calculate the average IMU bias for each sequence, addressing the issue of the lack of ground truth in public datasets. The generated average IMU biases are treated as training labels and successfully validated as reliable resources for improving the localization accuracy. The average IMU bias generation process and related data will be released to benefit the community.
@@
(iii) Extensive experiments on public and self-collected multi-platform datasets (handheld and autonomous vehicle) demonstrate significant improvements in localization precision and robustness.
The remainder of this paper is organized as follows. Section 2 reviews related studies on VIO and learning-based inertial estimation. Section 3 introduces the proposed IPNet framework in detail, including its network design, bias prior factor formulation, and the iterative label generation method. Section 4 presents the experimental setup, datasets, evaluation metrics, and quantitative results compared to baseline and state-of-the-art methods. Section 5 discusses the results, including inference speed, generalization ability, and the requirement of Vicon. Finally, Section 6 concludes the paper and discusses future research directions.
@@ II-A Visual inertial odometry
The VIO algorithm was initially developed using the Extended Kalman Filter (EKF), primarily due to the computational limitations of early systems. This line of research has continued to evolve, with notable contributions including ROVIO [11], MSCKF [12], OpenVINS [13], and SchurVINS [14]. With advancements in nonlinear optimization theory and improvements in computational power, optimization-based VIO methods have become the dominant approach. These methods integrate observational data and system constraints by formulating optimization problems, effectively overcoming the limitations of traditional filtering methods in handling complex nonlinear systems. Notable examples include OKVIS [15], VINS-Mono [16], ORB-SLAM3 [17], BASALT [18], Kimera [19], and DM-VIO [20].
However, existing methods often struggle to perform reliably under extreme conditions, such as low-texture environments or fast motion. These challenging scenarios can significantly degrade the accuracy of state estimation or even lead to system failures. Therefore, enhancing the robustness of VIO algorithms under degraded visual conditions remains a fundamental and ongoing challenge in the field.
@@ II-B Learning based inertial odometry
In addition to trajectory estimation, many learning-based methods focus on IMU noise reduction and uncertainty estimation. For example, Brossard et al. [26] employed Convolutional Neural Networks (CNN) for gyro denoising. Steinbrener et al. [27] compared different architectures for denoising IMU measurements and found that Long Short-Term Memory Networks (LSTM) outperform Transformer architectures, especially under varying data rates. AirIMU [28] introduced an encoder-decoder network structure to process raw IMU data, estimating IMU states and uncertainties, and outputting corrected acceleration and angular velocity values along with their associated uncertainties. Zeinali et al. [29] proposed MobileResNet, a CNN-based lightweight block that utilizes deep and pointwise convolutions to reduce computational costs and noise interference, thereby improving navigation and localization. Moreover, Yang et al. [30] integrated the Left-Invariant Extended Kalman Filter with a statistical neural network, proposing a learning-based dead reckoning algorithm that relies solely on inertial measurements and exhibits robust performance under sudden changes in lighting conditions.
However, existing methods still face challenges due to integration drift caused by accumulated noise and bias, which reduces the long-term accuracy of trajectory estimation. To address this issue, researchers have integrated IMU data with visual sensors and leverage deep learning to further improve both accuracy and robustness.
@@ II-C Learning based visual inertial odometry
Unlike existing methods, we restructured the IMU bias prediction paradigm and proposed an end-to-end average bias regression network (see Section III-B for details). Based on this, we designed a regularization prior factor that can be easily integrated into any VIO method.
Figure 3: The architecture of IPNet consists of encoder block, sequence modeling block and decoder block. The encoder block is mainly composed of four residual blocks and max-pooling layers, used to extract features and reduce the data dimension. The sequence modeling block is designed to further capture temporal dependencies, and the decoder block is responsible for performing regression to predict the bias of both acceleration and angular velocity.
@@ III Methodology
The overview of our proposed framework is shown in Fig. 2. We use VINS-Fusion [5] as the baseline and integrate the plug-and-play learning-based IMU bias prior factor into it. This integration addresses a key challenge in VIO frameworks: the degradation of localization precision and robustness caused by inaccurate bias estimation due to visual tracking errors. Furthermore, to facilitate the acquisition of IMU bias labels for network training, we introduce a method for iteratively solving the bias using the ground truth pose and velocity.
@@ III-A2 Preintegration
We follow the discrete expressions in [33] and express the displacement, velocity, and rotational preintegration terms as $\boldsymbol{\alpha}_{b_{k+1}}^{b_{k}}$$\boldsymbol{\beta}_{b_{k+1}}^{b_{k}}$$\gamma_{b_{k+1}}^{b_{k}}$.
$\boldsymbol{\alpha}_{b_{k+1}}^{b_{k}}$ $=\iint_{t\in\left[t_{k},t_{k+1}\right]}\mathbf{R}_{t}^{b_{k}}\left(\hat{\mathbf{a}}_{t}-\mathbf{b}_{a_{t}}-\mathbf{n}_{a}\right)dt^{2},$ (2)
$\boldsymbol{\beta}_{b_{k+1}}^{b_{k}}$ $=\int_{t\in\left[t_{k},t_{k+1}\right]}\mathbf{R}_{t}^{b_{k}}\left(\hat{\mathbf{a}}_{t}-\mathbf{b}_{a_{t}}-\mathbf{n}_{a}\right)dt,$
$\boldsymbol{\gamma}_{b_{k+1}}^{b_{k}}$ $=\int_{t\in\left[t_{k},t_{k+1}\right]}\frac{1}{2}\boldsymbol{\Omega}\left(\hat{\boldsymbol{\omega}}_{t}-\mathbf{b}_{w_{t}}-\mathbf{n}_{w}\right)\boldsymbol{\gamma}_{t}^{b_{k}}dt.$
@@ III-B IPNet model
In this work, we propose IPNet to directly predict the bias from raw IMU data according to the motion prior learned from existing trajectories [7]. The network consists of three blocks: encoder block, sequence modeling block, and decoder block, as shown in Fig.3. The raw IMU data represent the acceleration and angular velocity measurements in three dimensions, and we take a sample window of length $s$ as the input to the model, so that the measurements can be denoted as $\hat{\mathbf{a}}_{t}\in\mathbb{R}^{s\times 3}$ and $\hat{\boldsymbol{\omega}}_{t}\in\mathbb{R}^{s\times 3}$. We predict $n$ values for each window to avoid prediction fluctuations. Therefore, the model outputs the predicted bias $\hat{\mathbf{b}}_{a}\in\mathbb{R}^{n\times 3}$ and $\hat{\mathbf{b}}_{w}\in\mathbb{R}^{n\times 3}$. In this paper, we set $s$ to 1000 and $n$ to 50.
@@ III-B1 Encoder block
The encoder block mainly consists of several transformation layers to reshape the tensor and an encoder to capture local features and reduce data dimension. Firstly, concatenate $\hat{\mathbf{a}}_{t}$ and $\hat{\boldsymbol{\omega}}_{t}$ along the last dimension, and then use the transpose operation to swap the positions of the last two dimensions, so as to facilitate the subsequent convolution operations. The encoder contains four sets of residual blocks and max-pooling layers. The main path of residual block primarily utilizes 1D convolution, batch normalization, and PReLU activation, and the residual connection employs a pointwise convolution layer to combine the features of different channels and make the input and output have the same dimension, which are then added to the main path. Finally, a transpose operation is applied to restore the original dimensional order.
@@ III-B3 Decoder block
To accommodate the characteristics of acceleration bias and angular velocity bias, two parallel decoders predict accelerometer and gyroscope biases. Additionally, to avoid instability caused by fluctuations in model predictions, each decoder outputs $n$ predicted values, and their mean is used as the final prediction for each data window.
@@ III-B4 Loss function
Figure 4: IMU Bias Calculation Accuracy Experiment. The horizontal axis represents the mean of the ground truth bias, and the vertical axis represents the bias label calculated by our method. Ideally, they should be equal and distributed along a straight line with a slope of 1 and passing through the origin. The figure displays the bias distribution and fitting of different datasets. The gray area represents the confidence interval, used to characterize the 95% confidence interval of the fitting curve, while the yellow area represents the prediction interval, used to characterize the estimated range of future observation values under the 95% confidence level.
III-C IMU bias prior factor using IPNet
In this work, we integrate IPNet into any VIO system by forming an IMU bias factor using the IPNet. The main purpose of this factor is to provide a bounded value for the bias in the optimization procedure within the backend of VIO. It evaluates the deviation between the actual bias value and the known initial value in the form of residual, and incorporates this constraint into the optimization problem, helping to reduce state drift. The definition of its residual is:
@@ III-C IMU bias prior factor using IPNet
III-D IMU bias estimation for training label
A fundamental challenge in learning IMU bias priors is the absence of ground-truth bias values in most visual-inertial datasets. To address this, we propose an optimization-based method to compute per-sequence average biases, which serve as the supervision signal for IPNet. This design choice is motivated by a key observation: while IMU biases theoretically undergo random walk, their variation over a typical VIO sequence is often dominated by a relatively stable offset. Our solution proceeds as follows: first, we leverage the pose ground truth and raw IMU measurements to iteratively solve for the IMU bias that best aligns the pre-integrated trajectory with the ground truth. Then, we average these biases over the entire sequence to obtain a single, representative label. By regressing towards this sequence average, IPNet learns to predict a stable and representative bias prior that captures the dominant error component induced by the specific platform’s motion pattern, rather than attempting to track the noisy instantaneous bias. This approach effectively smoothes out the high-frequency noise inherent in the optimization-based inverse calculation and provides a robust learning target for the network.
The steps of the algorithm are as follows: First, the raw IMU data is pre-integrated to obtain the measured values of the pre-integrated terms. These values are then subtracted from the ground truth of the pre-integrated terms to calculate the error. Finally, $\mathbf{b}_{a_{t}}$ and $\mathbf{b}_{w_{t}}$ are inversely solved through iterative calculations based on the first-order term of the pre-integrated with respect to bias changes. The formula is as follows: